### v2

Abstract
We consider optimal control problems involving nonlinear ordinary differential equations with uncertain inputs. Using the sample average approximation, we obtain optimal control problems with ensembles of deterministic dynamical systems. Leveraging techniques for metric entropy bounds, we derive non-asymptotic Monte Carlo-type convergence rates for the ensemble-based solutions. Our theoretical framework is validated through numerical simulations on a harmonic oscillator problem and a vaccination scheduling problem for epidemic control under model parameter uncertainty.
keywords:
stochastic optimization, sample average approximation, Monte Carlo sampling, sample complexity, uncertain dynamical systems, optimal control
AMS subject classifications. 90C15, 90C30, 49K45, 49N60
1 Introduction
We consider the optimal control problem
$\min_{u\in L^{2}(0,1;\mathbb{R}^{m})}\,\mathbb{E}[F(x^{u}(1,\xi),\xi)]+\psi(u),$ (1)
where for each parameter $\xi\in\Xi$ and control $u(\cdot)\in L^{2}(0,1;\mathbb{R}^{m})$, $x^{u}(\cdot,\xi)=x(\cdot,\xi)$ solves the parameterized affine-control dynamical system
$\dot{x}(t,\xi)=f_{0}(x(t,\xi),\xi)+f_{1}(x(t,\xi),\xi)u(t)\quad\text{for a.e. }t\in(0,1),\quad x(0,\xi)=x_{0}(\xi),$ (2)
where $\Xi$ is a complete separable metric space equipped with its Borel sigma-algebra, $F\colon\mathbb{R}^{n}\times\Xi\to[0,\infty)$, $f_{0}\colon\mathbb{R}^{n}\times\Xi\to\mathbb{R}^{n}$ and $f_{1}\colon\mathbb{R}^{n}\times\Xi\to\mathbb{R}^{n\times m}$ are Carathéodory mappings, and $x_{0}\colon\Xi\to\mathbb{R}^{n}$ is measurable. The function $\psi\colon L^{2}(0,1;\mathbb{R}^{m})\to(-\infty,\infty]$ is proper, lower semicontinuous, and strongly convex with parameter $\alpha>0$. The parameterized initial value problem in (2) models uncertain right-hand sides and initial values. With some abuse of notation, we use $\xi$ to denote elements of $\Xi$ and a random element taking values in $\Xi$.
We use the sample average approximation (SAA) approach [10] to approximate the infinite dimensional optimization problem (1). Throughout the text, let $\xi^{1},\xi^{2},\ldots$ be independent and identically distributed $\Xi$-valued random elements defined on a complete probability space $(\Omega,\mathcal{F},P)$ such that each $\xi^{i}$ has the same distribution as $\xi$. We obtain the SAA problem
$\min_{u\in L^{2}(0,1;\mathbb{R}^{m})}\,\frac{1}{N}\sum_{i=1}^{N}F(x^{u}(1,\xi^{i}(\omega)),\xi^{i}(\omega))+\psi(u).$ (3)
The optimization problem (3) is an optimal control problem with an ensemble of $N$ dynamical systems. For the numerical solution of (3), we generate a realization of $\xi^{1}(\omega),\xi^{2}(\omega),\ldots,\xi^{N}(\omega)$ and solve (3) for these fixed elements. In the remainder of the text, we often omit explicit evaluations of the random elements $\xi^{i}$ at $\omega\in\Omega$.
1.1 Contributions
In this section, we summarize our four main contributions. Our summary uses standard notation, which is formally introduced in section 2. We define the parameterized integrand
$T(u,\xi)\coloneqq F(x^{u}(1,\xi),\xi).$
Furthermore, we define $\psi_{\alpha}(u)\coloneqq\psi(u)-(\alpha/2)\|u\|_{L^{2}(0,1;\mathbb{R}^{m})}^{2}$,
$g(u)\coloneqq\mathbb{E}[T(u,\xi)],\quad\text{and}\quad\widehat{g}_{N}(u)\coloneqq\frac{1}{N}\sum_{i=1}^{N}T(u,\xi^{i}).$ (4)
The following items list our four main contributions.
  1. (a)
    We establish nonasymptotic mean convergence rates for the SAA optimal values. Specifically, we show that for all $N\in\mathbb{N}$,
    $\mathbb{E}[|\widehat{v}_{N}^{*}-v^{*}|]\leq\frac{\text{Const}}{\sqrt{N}}\bigg(1+\frac{1}{\sqrt{\alpha}}\bigg),$ (5)
    where $v^{*}$ is the optimal value of (1) and $\widehat{v}_{N}^{*}$ is that of (3). Moreover, Const is a constant that does not depend on the sample size $N$ nor on the strong convexity parameter $\alpha$. However, it can depend on other problem data, such as the control’s dimension $m$.
  2. (b)
    We demonstrate nonasymptotic mean convergence rates for a criticality measure for (1) evaluated at SAA critical points: for each critical point $u_{N}^{*}\in\mathrm{dom}(\psi)$ of (3), that is, $\widehat{\chi}_{N}(u_{N}^{*})=0$, we show that for all $N\in\mathbb{N}$,
    $\mathbb{E}[\chi(u_{N}^{*})]\leq\frac{\text{Const}}{\sqrt{N}}\bigg(1+\frac{1}{\sqrt{\alpha}}\bigg),$ (6)
    where $\mathrm{Const}$ is as in (5) and the criticality measures $\chi$ and $\widehat{\chi}_{N}$ are defined by
    $\chi(u)\coloneqq\|u-\mathrm{prox}_{\psi_{\alpha}}(u-\nabla g(u)-\alpha u)\|_{L^{2}(0,1;\mathbb{R}^{m})}$ (7)
    and
    $\widehat{\chi}_{N}(u)\coloneqq\|u-\mathrm{prox}_{\psi_{\alpha}}(u-\nabla\widehat{g}_{N}(u)-\alpha u)\|_{L^{2}(0,1;\mathbb{R}^{m})}.$
  3. (c)
    Our contributions in (a) and (b) are derived using uniform expectation bounds for Hilbert space-valued “sub-Gaussian-type” Carathéodory mappings. Since this result may be of independent interest, we summarize it here. Let $H$ be a real, separable Hilbert space, let $C\subset H$ be a nonempty, closed set with diameter $D>0$, and let $G\colon C\times\Xi\to H$ be a Carathéodory mapping such that $\xi\mapsto G(x,\xi)$ is integrable and $\mathbb{E}[G(x,\xi)]=0$ for each $x\in C$. Furthermore, suppose that there exists a constant $M>0$ such that the sub-Gaussian-type estimate
    $\mathbb{E}[\cosh(\lambda\|G(x,\xi)-G(y,\xi)\|_{H})]\leq\exp(M^{2}\lambda^{2}\|x-y\|_{H}^{2}/2)\;\text{for all}\;x,y\in C,\;\lambda\geq 0$
    holds. Under these technical conditions, we establish the uniform expectation bound
    $\mathbb{E}[\sup_{x\in C}\,\|\widehat{G}_{N}(x)\|_{H}]\leq\mathbb{E}[\|\widehat{G}_{N}(x_{0})\|_{H}]+\frac{4\sqrt{3}M}{\sqrt{N}}\int_{0}^{D/2}\sqrt{\ln(2\mathcal{N}(C,\varepsilon))}\,\mathrm{d}\varepsilon,$ (8)
    where $\widehat{G}_{N}(x)\coloneqq(1/N)\sum_{i=1}^{N}G(x,\xi^{i})$, and $\mathcal{N}(C,\varepsilon)$ is the $\varepsilon$-covering number of $C$ (see section 2 for a definition). This uniform expectation bound is valid for each fixed $x_{0}\in C$.
  4. (d)
    Contributions (a) through (c) are technical in nature. Our final contribution involves formulating vaccination scheduling for epidemic control with random inputs as a risk-neutral optimal control problem. We validate our theoretical findings for this vaccination scheduling problem through numerical simulations.
1.2 Related work
The SAA approach has been used for various optimal control problems with uncertain dynamics, notably in optimizing batch reactors with model parameter uncertainty [26, 29]. Recent texts have established some of its theoretical properties, including asymptotic consistency and sample complexity. The paper [21] analyzes the almost sure epiconvergence of the SAA objective function and consistency statements for SAA critical points. Further asymptotic properties are demonstrated in [27]. Moreover, the work [16] establishes the optimal sample complexity of strongly convex stochastic optimization in Hilbert space.
Additionally, theoretical analyses of the SAA framework have been extended to other infinite dimensional stochastic nonconvex optimization problems with differential equation constraints. This includes the derivation of sample size estimates for semilinear PDE-constrained optimization under uncertainty [20] and the consistency analysis of sample-based approximations to risk-averse stochastic optimization problem with infinite dimensional decision spaces [19]. For risk-neutral linear elliptic PDE-constrained optimization, mean convergence rates for SAA optimal values and solutions, as well as a central limit theorem, are established in [24]. Parts of our theoretical analysis is based on the construction of a deterministic compact set containing the SAA critical points. This construction is inspired by the analysis developed in [20], but we use different statistical tools to derive convergence rates.
The right-hand side of (8) resembles the classical metric entropy bound for the suprema of sub-Gaussian processes (see, for example, [6, Thm. 3.1 on p. 95]). Metric entropy bounds are one of the tools for analyzing the SAA approach; see, for example, [14].
1.3 Manuscript overview
We present notation and collect basic terminology in section 2. In section 3, we formulate technical assumptions on the optimal control problem and the initial value problem. These assumptions enable us to establish the gradient and control regularity in sections 4 and 5. Our main results on the convergence rates for the SAA optimal values and critical points are presented in sections 6 and 7. These theoretical results are empirically validated in section 8 through two examples. Our appendices collect and establish results necessary for establishing the SAA convergence rates. While appendix A states essentially known first-order optimality conditions, appendix B derives bounds on the covering numbers of square-integrable mappings with square-integrable derivatives. Finally, appendix C concludes by establishing the uniform exponential tail bounds.
2 Notation
We collect basic notation and terminology. If not stated otherwise, equations involving random elements hold with probability one. Let $H$ be a Hilbert space with inner product $(\cdot,\cdot)_{H}$. We use $\|\cdot\|_{H}$ to denote the norm induced by its inner product $(\cdot,\cdot)_{H}$. For a Banach space $X$, $\|\cdot\|_{X}$ denotes its norm. For Banach spaces $X_{1}$, $\ldots$, $X_{m}$, we equip the Cartesian product $X_{1}\times\cdots\times X_{m}$ with the norm $\|(x_{1},\ldots,x_{m})\|_{X_{1}\times\cdots\times X_{m}}\coloneqq(\sum_{i=1}^{m}\|x_{i}\|_{X_{i}}^{2})^{1/2}$ if $X_{1}$, $\ldots$, $X_{m}$ are Hilbert spaces, and $\|(x_{1},\ldots,x_{m})\|_{X_{1}\times\cdots\times X_{m}}\coloneqq\sum_{i=1}^{m}\|x_{i}\|_{X_{i}}$ otherwise. We denote by $\mathbb{B}_{X}(0;r)$ the open unit norm ball with radius $r$ and by $\overline{\mathbb{B}}_{X}(0;r)$ its closure. If not stated otherwise, we equip $\mathbb{R}^{m}$ with the Euclidean norm $\|\cdot\|_{2}$. We denote the spectral norm of a matrix $A$ and more generally the operator norm of a linear operator $A$ between Euclidean spaces by $\|A\|_{2}$. We call an extended real-valued function $G$ defined on a Hilbert space $H$ strongly convex with parameter $\alpha>0$ if $G(\cdot)-(\alpha/2)\|\cdot\|_{H}^{2}$ is convex. The domain of a function $\phi$ is denoted by $\mathrm{dom}(\phi)$. For a proper, lower semicontinuous, convex function $\phi$ on a Hilbert space $H$, its proximity operator is defined by $\mathrm{prox}_{\phi}(x)\coloneqq\mathrm{argmin}_{y\in H}\,(1/2)\|x-y\|_{H}^{2}+\phi(y)$.
The spaces $L^{p}(0,1;\mathbb{R}^{m})$ $(1\leq p\leq\infty)$ denote the “usual” Lebesgue spaces equipped with their canonical norms. For $1\leq p\leq\infty$, we define the Sobolev space
$W_{1}^{p}(0,1;\mathbb{R}^{m})\coloneqq\{\,v\in L^{p}(0,1;\mathbb{R}^{m})\colon\,\dot{v}\in L^{p}(0,1;\mathbb{R}^{m})\,\}.$
Here, $\dot{v}$ denotes the distributional (time) derivative of $v$. While we generally use the “dot notation,” we occasionally write $D_{t}v$ in contexts where $\dot{v}$ is typographically inconvenient. We equip $W_{1}^{p}(0,1;\mathbb{R}^{m})$ with
$\|v\|_{W_{1}^{p}(0,1;\mathbb{R}^{m})}\coloneqq\big(\|v\|_{L^{p}(0,1;\mathbb{R}^{m})}^{p}+\|\dot{v}\|_{L^{p}(0,1;\mathbb{R}^{m})}^{p}\big)^{1/p}$
if $1\leq p<\infty$ and
$\|v\|_{W_{1}^{\infty}(0,1;\mathbb{R}^{m})}\coloneqq\|v\|_{L^{\infty}(0,1;\mathbb{R}^{m})}+\|\dot{v}\|_{L^{\infty}(0,1;\mathbb{R}^{m})}$
if $p=\infty$.
Let $H_{2}\subset H_{1}$ be convex sets in a real Hilbert space $H$. Adapting Definition 5.1.26 in [23], we say that $G\colon H_{1}\to\mathbb{R}$ is Gateaux differentiable relative to $H_{1}$ at a point $v\in H_{2}$ if there exists a linear bounded operator $DG(v)$ on $H$ such that for each $h\in H$ with $v+\lambda h\in H_{1}$ for some $\lambda>0$, $G$ is directionally differentiable at $v$ in the direction $h$ with directional derivative $DG(v)h$. We say that $G$ is Gateaux differentiable on $H_{2}$ relative to $H_{1}$ if it is Gateaux differentiable relative to $H_{1}$ at each $v\in H_{2}$. We denote by $\nabla G(v)$ the Riesz representation of $DG(v)$. (Different from Definition 5.1.26 in [23], we do not require $G$ be continuous). We denote the Gateaux derivative of a function $G$ by $DG$. The partial derivative of $G$ with respect to $v$ is denoted by $D_{v}G$.
For a metric space $(Y,d_{Y})$ and a nonempty, totally bounded subset $Y_{0}\subset Y$, we define the $\nu$-covering number $\mathcal{N}(Y_{0},Y,\nu)$ as the minimal number of points in a $\nu$-net of $Y_{0}$ in $Y$. We note that the elements in $\nu$-nets of $Y_{0}$ in $Y$ are not required to be contained in $Y_{0}$. Since $Y_{0}$ is totally bounded, $(Y_{0},d_{Y})$ is a totally bounded metric space. For abbreviation, we define $\mathcal{N}(Y_{0},\nu)\coloneqq\mathcal{N}(Y_{0},Y_{0};\nu)$.
3 Assumptions on optimal control problem
We state a number of assumptions on the optimal control problem (1) and the initial value problem (2), in addition to those stated in section 1. The assumptions ensure, for example, well-posedness of the optimal control problems, differentiability of sample average and expectation functions, and existence and uniqueness of solutions to the parameterized initial value problem. Moreover, these conditions allow us to demonstrate that the integrand in (1) has square integrable time derivatives for each parameter $\xi$. These conditions are mainly inspired by those used in [21, 23]. After formulating the first assumption, we comment on three key differences of our problem formulation and that of the constrained risk-neutral control problem in [21].
Our analysis is based on the following conditions on the control regularizer.
Assumption 1.
There exists a constant $r_{\psi}\in(0,\infty)$ such that $\|u\|_{L^{\infty}(0,1;\mathbb{R}^{m})}\leq r_{\psi}$ for all $u\in\mathrm{dom}(\psi)$.
Assumption 1 ensures that the domain of $\psi$ is a bounded subset in $L^{\infty}(0,1;\mathbb{R}^{m})$. Assumption 1 is, for example, satisfied for the sparsity-promoting control regularizer
$\psi(u)=(\alpha/2)\|u\|_{L^{2}(0,1;\mathbb{R}^{m})}^{2}+\beta\|u\|_{L^{1}(0,1;\mathbb{R}^{m})}+\iota_{\{u\in L^{2}(0,1;\mathbb{R}^{m})\colon\|u\|_{L^{\infty}(0,1;\mathbb{R}^{m})}\leq 1\}}(u),$
where $\alpha>0$, and $\beta\geq 0$. Moreover, $\iota_{Y}\colon L^{2}(0,1;\mathbb{R}^{m})\to\{0,\infty\}$ is the indicator function of the set $Y\subset L^{2}(0,1;\mathbb{R}^{m})$, that is, $\iota_{Y}(y)=0$ if $y\in Y$ and $\iota_{Y}(y)=\infty$ otherwise.
Our formulation of the control problem (1) differs from that of the constrained risk-neutral control problem in [21] in three key aspects: (i) we consider affine-linear controls in (2) instead of potentially nonaffine control dependence, (ii) we impose the strong convexity of $\psi$, and (iii) we omit the inclusion of additional $n$-dimensional decision variables that are added to the initial value $x_{0}(\xi)$ in (2). The first two aspects enable us to construct a deterministic compact set containing the SAA solutions. By omitting the additional, finite dimensional decision variables, we can present a concise sample complexity analysis focused on the infinite dimensional controls. Nonetheless, the inclusion of finite dimensional spaces remains a feasible direction for future work. Despite these modifications, we are still able to leverage the theoretical results from [21], including differentiability statements and Lipschitz constant estimates of the integrand.
The next assumption is adapted from Assumption 1 in [21]. It requires that the parameterized initial value problem has a solution and that the state be contained in a deterministic compact set.
Assumption 2.
There exists a convex, compact set $X_{0}\subset\mathbb{R}^{n}$ with $0\in X_{0}$ such that for each $\xi\in\Xi$ and $u\in\mathbb{B}_{L^{\infty}(0,1;\mathbb{R}^{m})}(0;2r_{\psi})$, it holds that $x^{u}(t,\xi)\in X_{0}$ for a.e. $t\in[0,1]$.
Assumption 2 implies that the parameterized initial value is bounded. More precisely, Assumption 2 ensures the existence of a constant $V_{0}\in(0,\infty)$ such that $\|x_{0}(\xi)\|_{2}\leq V_{0}$ for all $\xi\in\Xi$.
The next assumption is motivated by Assumption 5.6.2 in [23] and Assumptions 2, 3, and 5 in [21]. We define the dynamical system’s right-hand side $f\colon\mathbb{R}^{n}\times\mathbb{R}^{m}\times\Xi\to\mathbb{R}^{n}$ by
$f(x,u,\xi)\coloneqq f_{0}(x,\xi)+f_{1}(x,\xi)u.$ (9)
Assumption 3.
Let $r_{\psi}$ be as in Assumption 1, and let $X_{0}$ be as in Assumption 2.
  1. (a)
    For each $\xi\in\Xi$, $f_{0}(\cdot,\xi)$ and $f_{1}(\cdot,\xi)$ are continuously differentiable, and for each $x\in X_{0}$, $D_{x}f_{0}(x,\cdot)$, and $D_{x}f_{1}(x,\cdot)$ are measurable.
  2. (b)
    There exists a constant $L_{f}\in(0,\infty)$ such that for each $\xi\in\Xi$, $f_{0}(\cdot,\xi)$, $f_{1}(\cdot,\xi)$, $D_{x}f_{0}(\cdot,\xi)$, and $D_{x}f_{1}(\cdot,\xi)$ are Lipschitz continuous on an open neighborhood of $X_{0}$ with Lipschitz constant $L_{f}$. Moreover, $\|f_{0}(0,\xi)\|_{2}\leq L_{f}$ for all $\xi\in\Xi$, and $\|f_{1}(x,\xi)\|_{2}\leq L_{f}$ and $\|D_{x}f_{1}(x,\xi)\|_{2}\leq L_{f}$ for all $(x,\xi)\in X_{0}\times\Xi$.
  3. (c)
    For each $\xi\in\Xi$, $F(\cdot,\xi)$ is continuously differentiable, and for each $x\in X_{0}$, $F(x,\cdot)$ and $D_{x}F(x,\cdot)$ are measurable.
  4. (d)
    There exists a constant $L_{F}^{\prime}\in(0,\infty)$ such that for each $\xi\in\Xi$, $F(\cdot,\xi)$ and $D_{x}F(\cdot,\xi)$ are Lipschitz continuous on an open neighborhood of $X_{0}$ with Lipschitz constant $L_{F}^{\prime}$.
Assumptions 1, 3 and 2 ensure that there exists a neighborhood $\tilde{X}_{0}$ of $X_{0}$, such that for each $\xi\in\Xi$, $f(\cdot,\cdot,\xi)$ and $D_{x}f(\cdot,\cdot,\xi)$ are Lipschitz continuous on $\tilde{X}_{0}\times\{u\in\mathbb{R}^{m}\colon\|u\|_{2}\leq 2r_{\psi}\}$ with Lipschitz constant
$L_{f}^{\prime}\coloneqq L_{f}\sqrt{(1+2r_{\psi})^{2}+1}.$
Hence for all $x\in X_{0}$, $\xi\in\Xi$, and $u\in\mathbb{R}^{m}$ with $\|u\|_{2}\leq 2r_{\psi}$,
$\begin{aligned} \|f(x,u,\xi)\|_{2}&\leq L_{f}^{\prime}(\|x\|_{2}+2r_{\psi}+1),\quad\text{and}\quad&\|D_{x}f(x,u,\xi)\|_{2}&\leq L_{f}^{\prime}.\end{aligned}$ (10)
We conclude this section by discussing Assumption 2.
Remark 3.1.
For globally Lipschitz continuous right-hand sides $f(\cdot,\cdot,\xi)$ with Lipschitz constants independent of $\xi$, standard results, such as Proposition 5.6.5 in [23], can be used to construct a set $X_{0}$ satisfying Assumption 2. Following the approach in [21], we assume the existence of such a set $X_{0}$. This enables us to work with right-hand sides that are Lipschitz continuous only on the compact set $X_{0}$, allowing a broader class of problems to be considered.
4 Gradient regularity via adjoint and state regularity
The section’s main purpose is to show that the objective function’s gradient is contained in a compact subset of the feasible set. The analysis performed is based on the results provided in [21, 23].
We recall that the control problem’s integrand $T\colon\mathbb{B}_{L^{\infty}(0,1;\mathbb{R}^{m})}(0;2r_{\psi})\times\Xi\to[0,\infty)$ is defined by
$T(u,\xi)\coloneqq F(x^{u}(1,\xi),\xi).$
If Assumptions 1, 3 and 2 hold and $\xi\in\Xi$, then Corollary 5.6.9 in [23] ensures that $T(\cdot,\xi)$ is Gateaux differentiable on $\mathrm{dom}(\psi)$ relative to $\mathbb{B}_{L^{\infty}(0,1;\mathbb{R}^{m})}(0;2r_{\psi})$ with
$[\nabla_{u}T(u,\xi)](t)=f_{1}(x^{u}(t,\xi),\xi)^{T}p^{u}(t,\xi)\quad\text{for a.e. }t\in(0,1),$ (11)
where $p^{u}(\cdot,\xi)=p(\cdot,\xi)$ solves for a.e. $t\in(0,1)$ the parameterized adjoint equation
$\dot{p}(t,\xi)=-D_{x}f(x^{u}(t,\xi),u(t),\xi)^{T}p(t,\xi),\quad p(1,\xi)=\nabla_{x}F(x^{u}(1,\xi),\xi).$ (12)
Building on the initial discussion, the section’s main purpose is to show that $\nabla_{u}T(u,\xi)$ is contained in a fixed deterministic ball in $W_{1}^{2}(0,1;\mathbb{R}^{m})$ for each $(u,\xi)\in\mathrm{dom}(\psi)\times\Xi$. This result is formulated in the following theorem, with an explicit expression for the ball’s radius.
Theorem 4.1.
If Assumptions 1, 3 and 2 hold, then for each $(u,\xi)\in\mathrm{dom}(\psi)\times\Xi$,
$\nabla_{u}T(u,\xi)\in\overline{\mathbb{B}}_{W_{1}^{\infty}(0,1;\mathbb{R}^{m})}(0;R)\subset\overline{\mathbb{B}}_{W_{1}^{2}(0,1;\mathbb{R}^{m})}(0;R),$
where
$R\coloneqq R_{0}+L_{f}^{\prime}\big[(1+2r_{\psi}+V_{0})\mathrm{e}^{L_{f}^{\prime}}+2r_{\psi}+2\big]R_{0}\quad\text{with}\quad R_{0}$ $\coloneqq L_{f}^{\prime}(1+L_{F}^{\prime})\mathrm{e}^{L_{f}^{\prime}}.$ (13)
We prepare our proof of Theorem 4.1. The following lemma collects basic properties of $T$ mainly taken from [21]. We define
$L_{T}^{\prime}\coloneqq\sqrt{2}L_{F}^{\prime}(L_{f}^{\prime}+1)\mathrm{e}^{L_{f}^{\prime}+1}.$ (14)
Lemma 4.2.
If Assumptions 1, 3 and 2 hold, then the following statements are valid.
  1. (a)
    For each $\xi\in\Xi$, $T(\cdot,\xi)$ is Lipschitz continuous with Lipschitz constant $L_{T}^{\prime}$, and Gateaux differentiable on $\mathrm{dom}(\psi)$ relative to $\mathbb{B}_{L^{\infty}(0,1;\mathbb{R}^{m})}(0;2r_{\psi})$.
  2. (b)
    There exists a constant $L_{\nabla T}^{\prime}\in(0,\infty)$ such that $\nabla_{u}T(\cdot,\xi)$ is Lipschitz continuous on $\mathrm{dom}(\psi)$ with Lipschitz constant $L_{\nabla T}^{\prime}$ for each $\xi\in\Xi$.
  3. (c)
    The mappings $T$ and $\nabla_{u}T$ are Carathéodory mappings on $\mathrm{dom}(\psi)\times\Xi$.
  4. (d)
    The function $u\mapsto\mathbb{E}[T(u,\xi)]$ is Gateaux differentiable on $\mathrm{dom}(\psi)$ relative to $\mathbb{B}_{L^{\infty}(0,1;\mathbb{R}^{m})}(0;2r_{\psi})$.
Proof.
  1. (a)
    The Lipschitz statement follows from the proof of Lemma 5.6.7 in [23], since $L_{f}^{\prime}+1\geq 1$. See also Lemma 3.5 in [21]. Furthermore, Corollary 5.6.9 in [23] ensures the Gateaux differentiability statement.
  2. (b)
    This statement follows from an application of Lemma 5.2 in [21].
  3. (c)
    According to Lemma 3.5 in [21], $T$ is a Carathéodory mapping. Using Lemma 2.3 in [21], composition rules, and (11), we can show that $\nabla_{u}T(u,\cdot)$ is measurable on $\Xi$. See also Lemma 5.2 in [21].
  4. (d)
    An application of Lemma 4.3 in [21] yields the assertion.
Using chain and product rules (cf. [13, p. 340] and [1, Cor. 3.2]), and (11), we find that for a.e. $t\in(0,1)$,
$D_{t}[\nabla_{u}T(u,\xi)](t)=D_{t}[f_{1}(x^{u}(t,\xi),\xi)^{T}]p^{u}(t,\xi)+f_{1}(x^{u}(t,\xi),\xi)^{T}\dot{p}^{u}(t,\xi).$ (15)
This identity allows us to establish stability estimates on the gradient’s time derivative.
Lemma 4.3.
If Assumptions 1, 3 and 2 hold, then the following statements hold true.
  1. (a)
    For each $(u,\xi)\in\mathbb{B}_{L^{\infty}(0,1;\mathbb{R}^{m})}(0;2r_{\psi})\times\Xi$, the dynamical system (2) has a unique solution $x^{u}(\cdot,\xi)\in W_{1}^{\infty}(0,1;\mathbb{R}^{n})$, and for a.e. $t\in[0,1]$,
    $\begin{aligned} \|x^{u}(t,\xi)\|_{2}&\leq(1+2r_{\psi}+V_{0})\mathrm{e}^{L_{f}^{\prime}},\\ \|\dot{x}^{u}(t,\xi)\|_{2}&\leq L_{f}^{\prime}(\|x^{u}(t,\xi)\|_{2}+2r_{\psi}+1),\end{aligned}$ (16)
    where we recall that $V_{0}\in(0,\infty)$ is such that $\|x_{0}(\xi)\|_{2}\leq V_{0}$ for all $\xi\in\Xi$.
  2. (b)
    For each $(u,\xi)\in\mathbb{B}_{L^{\infty}(0,1;\mathbb{R}^{m})}(0;2r_{\psi})\times\Xi$, the adjoint system (12) has a unique solution $p^{u}(\cdot,\xi)\in W_{1}^{\infty}(0,1;\mathbb{R}^{n})$, and for a.e. $t\in[0,1]$,
    $\begin{aligned} \|p^{u}(t,\xi)\|_{2}&\leq(1+L_{F}^{\prime})\mathrm{e}^{L_{f}^{\prime}},\\ \|\dot{p}^{u}(t,\xi)\|_{2}&\leq L_{f}^{\prime}\|p^{u}(t,\xi)\|_{2}.\end{aligned}$ (17)
  3. (c)
    For all $(u,\xi)\in\mathrm{dom}(\psi)\times\Xi$,
    $\|\nabla_{u}T(u,\xi)\|_{L^{\infty}(0,1;\mathbb{R}^{m})}\leq L_{f}^{\prime}(1+L_{F}^{\prime})\mathrm{e}^{L_{f}^{\prime}}.$
  4. (d)
    For all $(u,\xi)\in\mathrm{dom}(\psi)\times\Xi$,
    $\|D_{t}[\nabla_{u}T(u,\xi)]\|_{L^{\infty}(0,1;\mathbb{R}^{m})}$ $\leq(L_{f}^{\prime})^{2}(1+L_{F}^{\prime})\mathrm{e}^{L_{f}^{\prime}}$
    $\quad\cdot\big[(1+2r_{\psi}+V_{0})\mathrm{e}^{L_{f}^{\prime}}+2r_{\psi}+2\big].$
Proof.
  1. (a)
    The existence and uniqueness follow from Proposition 5.6.5 in [23]. Using (10), we obtain for all $t\in[0,1]$,
    $\|x^{u}(t,\xi)\|_{2}\leq\|x_{0}(\xi)\|_{2}+L_{f}^{\prime}\int_{0}^{t}\big(\|x^{u}(\tau,\xi)\|_{2}+2r_{\psi}+1)\,\mathrm{d}\tau.$
    Defining $y(t)\coloneqq\|x^{u}(t,\xi)\|_{2}+2r_{\psi}+1$, this estimate ensures $y(t)\leq y(0)+L_{f}^{\prime}\int_{0}^{t}y(\tau)\,\mathrm{d}\tau$ for all $t\in[0,1]$. Now, applying Gronwall’s inequality yields the first stability estimate. Using (2), and (10), we obtain the second stability estimate.
  2. (b)
    The existence and uniqueness of the solution to the adjoint equation follows from Corollary 5.6.9 in [23].
    The mean value theorem ensures $\|\nabla_{x}F(x,\xi)\|_{2}\leq L_{F}^{\prime}$ for all $(x,\xi)\in X_{0}\times\Xi$. Combined with (10), and Exercise 5.6.6 in [23], we obtain the first adjoint state stability estimate. The second is a consequence of the adjoint equation (12).
  3. (c)
    Using (10), and (11), we have for all $(u,\xi)\in\mathrm{dom}(\psi)\times\Xi$ and a.e. $t\in(0,1)$,
    $\|[\nabla_{u}T(u,\xi)](t)\|_{2}\leq\|f_{1}(x^{u}(t,\xi),\xi)\|_{2}\|p^{u}(t,\xi)\|_{2}\leq L_{f}^{\prime}\|p^{u}(t,\xi)\|_{2}.$
    Combined with (17), we obtain the estimate.
  4. (d)
    Using (10) and (15), we obtain
    $\|D_{t}[\nabla_{u}T(u,\xi)]\|_{L^{\infty}(0,1;\mathbb{R}^{m})}$ $\leq L_{f}^{\prime}\|p^{u}(\cdot,\xi)\|_{L^{\infty}(0,1;\mathbb{R}^{n})}\|\dot{x}^{u}(\cdot,\xi)\|_{L^{\infty}(0,1;\mathbb{R}^{n})}$
    $\quad+L_{f}^{\prime}\|\dot{p}^{u}(\cdot,\xi)\|_{L^{\infty}(0,1;\mathbb{R}^{n})}.$
    Now, (16) and (17) ensure the bound.
The proof of Theorem 4.1 is now a direct consequence of Lemma 4.3.
Proof of Theorem 4.1.
Since
$\|\nabla_{u}T(u,\xi)\|_{W_{1}^{\infty}(0,1;\mathbb{R}^{m})}=\|\nabla_{u}T(u,\xi)\|_{L^{\infty}(0,1;\mathbb{R}^{m})}+\|D_{t}[\nabla_{u}T(u,\xi)]\|_{L^{\infty}(0,1;\mathbb{R}^{m})},$
Lemma 4.3 ensures the assertion. ∎
5 Control regularity via gradient regularity
We show that SAA critical points are contained in a compact deterministic subset in the control space $L^{2}(0,1;\mathbb{R}^{m})$. The technical results presented in this section form the basis for establishing our mean convergence rates in sections 6 and 7. We recall that $\chi$ is the true problem’s criticality measure, while $\widehat{\chi}_{N}$ is that of the SAA problem, as defined in the introduction. We call $u\in\mathrm{dom}(\psi)$ a critical point of (1) if $\chi(u)=0$. Moreover, we refer to $u_{N}\in\mathrm{dom}(\psi)$ as a critical point of the SAA problem if $\widehat{\chi}_{N}(u_{N})=0$.
We define the subset $\mathcal{C}\subset L^{2}(0,1;\mathbb{R}^{m})$ by
$\mathcal{C}\coloneqq\mathrm{prox}_{(1/\alpha)\psi_{\alpha}}(-(1/\alpha)\overline{\mathbb{B}}_{W_{1}^{2}(0,1;\mathbb{R}^{m})}(0;R)),$ (18)
where we recall $\psi_{\alpha}(u)=\psi(u)-(\alpha/2)\|u\|_{L^{2}(0,1;\mathbb{R}^{m})}^{2}$.
Our next theorem establishes the covering numbers of $\mathcal{C}$. Moreover, it demonstrates that each critical point of the control problem (1) and of its SAA problem (3) is contained in $\mathcal{C}$.
Theorem 5.1.
If Assumptions 1, 3 and 2 hold, then the following statements are valid.
  1. (a)
    The diameter of $\mathcal{C}$ in $L^{2}(0,1;\mathbb{R}^{m})$ is at most $2r_{\psi}$.
  2. (b)
    For all $\nu>0$, the $\nu$-covering number of $\mathcal{C}$ with respect to the $L^{2}(0,1;\mathbb{R}^{m})$-norm satisfies
    $\log_{2}(\mathcal{N}(\mathcal{C},\nu))\leq\varrho\sqrt{m}(Rm/(\alpha\nu)),$
    where $\varrho>0$ is a constant independent of $m$, $\nu$, $R$, and $\alpha$.
  3. (c)
    We have
    $\int_{0}^{r_{\psi}}\sqrt{\ln(2\mathcal{N}(\mathcal{C},\varepsilon))}\,\mathrm{d}\varepsilon\leq 2\sqrt{r_{\psi}\varrho(Rm^{3/2}/\alpha)}+r_{\psi}.$
  4. (d)
    Each critical point of (1) and of (3) is contained in $\mathcal{C}$.
Proof.
  1. (a)
    We have $\mathcal{C}\subset\mathrm{dom}(\psi_{\alpha})=\mathrm{dom}(\psi)$. Assumption 1 ensures that $\mathrm{dom}(\psi)\subset\overline{\mathbb{B}}_{L^{\infty}(0,1;\mathbb{R}^{m})}(0;r_{\psi})$. Hence the diameter of $\mathcal{C}$ in $L^{2}(0,1;\mathbb{R}^{m})$ is at most $2r_{\psi}$.
  2. (b)
    We define the constant $\varepsilon\coloneqq(\alpha/R)\nu$. Proposition B.1 ensures the existence of points $v_{1},\ldots,v_{K}\in\overline{\mathbb{B}}_{W_{1}^{2}(0,1;\mathbb{R}^{m})}(0;1)$ such that $\|v-v_{k(v)}\|_{L^{2}(0,1;\mathbb{R}^{m})}\leq\varepsilon$ for all $v\in\overline{\mathbb{B}}_{W_{1}^{2}(0,1;\mathbb{R}^{m})}(0;1)$, where $k(v)\coloneqq\mathrm{argmin}_{k\in\{1,\ldots,K\}}\|v-v_{k}\|_{L^{2}(0,1;\mathbb{R}^{m})}$. Additionally, $K\leq 2^{\varrho\sqrt{m}(m/\varepsilon)}$, where $\varrho>0$ independent of $m$ and $\varepsilon$.
    Let $w\in\overline{\mathbb{B}}_{W_{1}^{2}(0,1;\mathbb{R}^{m})}(0;R)$. Then $v\coloneqq w/R\in\overline{\mathbb{B}}_{W_{1}^{2}(0,1;\mathbb{R}^{m})}(0;1)$. Since proximity operators are nonexpansive,
    $\|\mathrm{prox}_{(1/\alpha)\psi_{\alpha}}(-(1/\alpha)w)-\mathrm{prox}_{(1/\alpha)\psi_{\alpha}}(-(1/\alpha)Rv_{k(v)})\|_{L^{2}(0,1;\mathbb{R}^{m})}$
    $\quad\leq(1/\alpha)\|w-Rv_{k(v)}\|_{L^{2}(0,1;\mathbb{R}^{m})}=(R/\alpha)\|v-v_{k(v)}\|_{L^{2}(0,1;\mathbb{R}^{m})}\leq(R/\alpha)\varepsilon=\nu.$
  3. (c)
    Combining the previous part and Example C.6 (applied with $D\leq 2r_{\psi}$ and $c=\varrho\sqrt{m}(Rm/\alpha)$), we obtain the integral bound.
  4. (d)
    Let $\bar{u}$ be a critical point of (1). Lemmas A.1 and 4.2 ensure
    $\bar{u}=\mathrm{prox}_{(1/\alpha)\psi_{\alpha}}(-(1/\alpha)\nabla_{u}\mathbb{E}[T(\bar{u},\xi)]).$
    Moreover, Theorem 4.1 and the dominated convergence theorem yield $\nabla_{u}\mathbb{E}[T(\bar{u},\xi)]=\mathbb{E}[\nabla_{u}T(\bar{u},\xi)]\in L^{2}(0,1;\mathbb{R}^{m})$. Theorem 4.1 also ensures $\nabla_{u}T(u,\xi)\in W_{1}^{2}(0,1;\mathbb{R}^{m})$ for all $(u,\xi)\in\mathrm{dom}(\psi)\times\Xi$. Since $W_{1}^{2}(0,1;\mathbb{R}^{m})$ is reflexive, $\overline{\mathbb{B}}_{W_{1}^{2}(0,1;\mathbb{R}^{m})}(0;R)$ is closed in $L^{2}(0,1;\mathbb{R}^{m})$. Combined with Proposition 1.2.12 in [9], we find that $\mathbb{E}[\nabla_{u}T(\bar{u},\xi)]\in W_{1}^{2}(0,1;\mathbb{R}^{m})$. Hence $\bar{u}\in\mathcal{C}$.
    Now, let $\bar{u}_{N}$ be a critical point of the SAA problem (3). Lemma A.1 yields $\bar{u}_{N}=\mathrm{prox}_{(1/\alpha)\psi_{\alpha}}(-(1/\alpha)\nabla\widehat{g}_{N}(\bar{u}_{N}))$. Since $\nabla_{u}T(u,\xi)\in W_{1}^{2}(0,1;\mathbb{R}^{m})$ for all $(u,\xi)\in\mathrm{dom}(\psi)\times\Xi$, we have $\nabla\widehat{g}_{N}(\bar{u}_{N})=(1/N)\sum_{i=1}^{N}\nabla_{u}T(\bar{u}_{N},\xi^{i})\in W_{1}^{2}(0,1;\mathbb{R}^{m})$. Hence $\bar{u}_{N}\in\mathcal{C}$.
6 Convergence rates of optimal values
We establish mean convergence rates for SAA optimal values, our first main contribution. Throughout the section we assume that the control problems (1) and (3) have solutions. The existence of solutions can be established using standard arguments, once it is shown that the parameterized objective function is weakly lower semicontinuous in its first argument; that is, $T(\cdot,\xi)$ is weakly lower semicontinuous on $\mathrm{dom}(\psi)$ for each $\xi\in\Xi$. This property follows by showing that the mapping $\mathrm{dom}(\psi)\ni u\mapsto x^{u}(1,\xi)\in\mathbb{R}^{n}$ is weakly continuous. Lemma 2.4 in [27] proves a weak-to-strong continuity result of solution operators for control-affine initial value problems. Its proof may be adapted to establish the weak continuity required here. Now, let $u^{*}$ be a solution to (1) and for each $N\in\mathbb{N}$, let $u_{N}^{*}\colon\Omega\to L^{2}(0,1;\mathbb{R}^{m})$ be measurable such that for each $\omega\in\Omega$, $u_{N}^{*}(\omega)$ is a solution to (3). Moreover, we denote by $v^{*}$ the optimal value of (1) and by $\widehat{v}_{N}^{*}\colon\Omega\to\mathbb{R}$ that of (3), which we also assume to be measurable. Along with the existence of solutions, the measurability of the SAA optimal values and solutions can be established using standard arguments.
Theorem 6.1.
If Assumptions 1, 3 and 2 hold and $u_{0}\in\mathcal{C}$, then for all $N\in\mathbb{N}$,
$\mathbb{E}[|\widehat{v}_{N}^{*}-v^{*}|]$ $\leq\frac{(\mathbb{E}[(T(u_{0},\xi)-\mathbb{E}[T(u_{0},\xi)])^{2}])^{1/2}}{\sqrt{N}}+\frac{16\sqrt{3}L_{T}^{\prime}}{\sqrt{N}}\Bigg[\bigg(\frac{r_{\psi}\varrho Rm^{3/2}}{\alpha}\bigg)^{1/2}+r_{\psi}\Bigg].$
Proof.
Since each solution to (1) is a critical point of (1), Theorem 5.1 ensures that the solution set of (1) is contained in $\mathcal{C}$. In particular, $u^{*}\in\mathcal{C}$. Analogously, each SAA solution is contained in $\mathcal{C}$. Hence,
$|\widehat{v}_{N}^{*}-v^{*}|\leq\sup_{u\in\mathcal{C}}\bigg|\mathbb{E}[T(u,\xi)]-\frac{1}{N}\sum_{i=1}^{N}T(u,\xi^{i})\bigg|.$
Next, we estimate the supremum’s expected value using Theorem C.1. For each $\xi\in\Xi$, Lemma 4.2 ensures that $T(\cdot,\xi)$ is Lipschitz continuous with Lipschitz constant $L_{T}^{\prime}$ on $\mathrm{dom}(\psi)$. We define $G(u,\xi)\coloneqq T(u,\xi)-\mathbb{E}[T(u,\xi)]$. For each $\xi\in\Xi$, $G(\cdot,\xi)$ is Lipschitz continuous with Lipschitz constant $2L_{T}^{\prime}$ on $\mathrm{dom}(\psi)$. Hence $G$ satisfies (21) with $M=2L_{T}^{\prime}$. Now Theorems C.1 and C.2 ensure
$\mathbb{E}[|\widehat{v}_{N}^{*}-v^{*}|]$ $\leq\frac{(\mathbb{E}[(T(u_{0},\xi)-\mathbb{E}[T(u_{0},\xi)])^{2}])^{1/2}}{\sqrt{N}}+\frac{8\sqrt{3}L_{T}^{\prime}}{\sqrt{N}}\int_{0}^{r_{\psi}}\sqrt{\ln(2\mathcal{N}(\mathcal{C},\varepsilon))}\,\mathrm{d}\varepsilon.$
Combined with the entropy integral bound provided by Theorem 5.1, we obtain the assertion. ∎
7 Convergence rates for criticality measures
We establish mean convergence rates for SAA critical points, our second main contribution. We measure accuracy of SAA critical points via the criticality measure defined in (7).
Theorem 7.1.
Let Assumptions 1, 3 and 2 hold. For each $N\in\mathbb{N}$, let the control $u_{N}^{*}\colon\Omega\to L^{2}(0,1;\mathbb{R}^{m})$ be measurable such that for each $\omega\in\Omega$, $u_{N}^{*}(\omega)\in\mathrm{dom}(\psi)$ is a critical point of the SAA problem (3). If $u_{0}\in\mathcal{C}$, then for all $N\in\mathbb{N}$,
$\mathbb{E}[\chi(u_{N}^{*})]$ $\leq\frac{(\mathbb{E}\|\nabla_{u}T(u_{0},\xi)-\mathbb{E}[\nabla_{u}T(u_{0},\xi)]\|_{L^{2}(0,1;\mathbb{R}^{m})}^{2})^{1/2}}{\sqrt{N}}$
$\quad+\frac{16\sqrt{3}L_{\nabla T}^{\prime}}{\sqrt{N}}\Bigg[\bigg(\frac{r_{\psi}\varrho Rm^{3/2}}{\alpha}\bigg)^{1/2}+r_{\psi}\Bigg].$
Proof.
Our proof follows steps similar to those in the proof of Theorem 6.1. For all $u\in\mathrm{dom}(\psi)$, the nonexpansiveness of proximity operators ensures
$|\chi(u)-\widehat{\chi}_{N}(u)|\leq\|\nabla g(u)-\nabla\widehat{g}_{N}(u)\|_{L^{2}(0,1;\mathbb{R}^{m})}.$
Combined with $\widehat{\chi}_{N}(u_{N}^{*})=0$ and $u_{N}^{*}\in\mathcal{C}$, we obtain
$\mathbb{E}[\chi(u_{N}^{*})]\leq\mathbb{E}\big[\sup_{u\in\mathcal{C}}\,\|\nabla g(u)-\nabla\widehat{g}_{N}(u)\|_{L^{2}(0,1;\mathbb{R}^{m})}\big].$
Next, we estimate the above right-hand side using Theorem C.1. For each $\xi\in\Xi$, Lemma 4.2 ensures that $\nabla_{u}T(\cdot,\xi)$ is Lipschitz continuous with Lipschitz constant $L_{\nabla T}^{\prime}$ on $\mathrm{dom}(\psi)$. We define $G(u,\xi)\coloneqq\nabla_{u}T(u,\xi)-\mathbb{E}[\nabla_{u}T(u,\xi)]$. For each $\xi\in\Xi$, $G(\cdot,\xi)$ is Lipschitz continuous with Lipschitz constant $2L_{\nabla T}^{\prime}$ on $\mathrm{dom}(\psi)$. Hence $G$ satisfies (21) with $M=2L_{\nabla T}^{\prime}$. Combined with Theorems 5.1, C.1, C.1 and C.2, we obtain the assertion. ∎
8 Numerical illustrations
We consider two numerical examples. The main purpose of our numerical illustrations is to demonstrate the typical Monte Carlo convergence rate for the SAA optimal values and SAA critical points established in sections 6 and 7. Before discussing these examples, we provide details on the discretization of the control problems, implementation details, and a description of our simulation output. We also introduce some terminology.
We refer to
$\min_{u\in L^{2}(0,1;\mathbb{R}^{m})}\,F(x^{u}(1,\mathbb{E}[\xi]),\mathbb{E}[\xi])+\psi(u),$
as the nominal problem to (1). It is obtained by replacing the random element $\xi$ by its expected value and simulating the dynamical system (2) with this expected value. For both numerical examples, we visualize the nominal solution and the solution to a certain SAA problem, which is introduced below and referred to as the reference problem. These graphical illustrations demonstrate that nominal solutions and reference solutions can have different characteristics.
To approximate the optimal value of (1) and the criticality measures $\chi$, we generated $N_{\text{ref}}\coloneqq 4096$ samples using a realization of a scrambled Sobol’ sequence and transformed these samples to take values in $\Xi$. For both examples, $\Xi$ is the Cartesian product of finitely many closed intervals and the components of $\xi$ are uniformly distributed. We denote the samples by $\zeta^{1},\ldots,\zeta^{N_{\text{ref}}}$, and formulate the reference problem
$\min_{u\in L^{2}(0,1;\mathbb{R}^{m})}\,\frac{1}{N_{\text{ref}}}\sum_{i=1}^{N_{\text{ref}}}F(x^{u}(1,\zeta^{i}),\zeta^{i})+\psi(u).$
Moreover, we approximate the criticality measure $\chi$ by
$\chi_{\text{ref}}(u)\coloneqq\Big\|u-\mathrm{prox}_{\psi_{\alpha}}\Big(u-\frac{1}{N_{\text{ref}}}\sum_{i=1}^{N_{\text{ref}}}\nabla_{u}T(u,\zeta^{i})-\alpha u\Big)\Big\|_{L^{2}(0,1;\mathbb{R}^{m})}.$
In addition to providing plots of the nominal and reference solutions, we compare their quality using performance metrics from [4]. We define the (reference) expected performance of the nominal policy (EPP) as the reference SAA objective evaluated at the nominal solution, and the reference performance (RP) as the optimal value of the reference SAA problem. The reference value of the stochastic solution (VSS) is defined as $\text{EPP}-\text{RP}$, and the relative VSS as $\text{VSS}/\text{EPP}$.
The dynamical systems are discretized using multiple shooting with a fixed step explicit fourth-order Runge–Kutta method. The control space is discretized using piecewise constant functions defined on a uniform grid over $0,t_{f}]$ with $q\coloneqq 50$ subintervals, where $t_{f}>0$ is a fixed final time. Consequently, the solutions to the SAA problems and the criticality measures depend on the discretization parameter $q$. For notational convenience, this discretization parameter is omitted in the notation. We generated Sobol’ sequences using SciPy’s quasi-Monte Carlo capabilities [25]. To model and solve the discretized SAA problems, we used the package EnsembleControl [17], which is built on the software framework CasADi [2], and the optimization solver IPOPT [31]. We used IPOPT with its default settings. Our simulation code and output are archived at [18]. Simulations were performed on a laptop equipped with a 12th Gen Intel(R) Core(TM) i7-1260P processor and 16 GB of RAM.
8.1 Harmonic oscillator
Figure 1: For the harmonic oscillator problem formulated in section 8.1, nominal solution (left) and reference solution (right).
Figure 2: For the harmonic oscillator problem formulated in section 8.1, convergence rate of the SAA optimal values (left) and of the reference criticality measure evaluated at SAA critical points (right). The empirical means $\widehat{\mathbb{E}}$ were computed using 5050 replications. The convergence rates were computed using least squares.
We consider a harmonic oscillator inspired by the problem formulation in [21, sect. 7.1]. Our optimization problem is given by
$\min_{u\in L^{2}(0,1;\mathbb{R}^{2})}\,(1/2)\mathbb{E}[x_{1}^{u}(1,\xi)^{2}+x_{2}^{u}(1,\xi)^{2}]+\psi(u),$
where for each parameter $\xi\in\Xi\subset\mathbb{R}^{5}$ and control $u(\cdot)\in L^{2}(0,1;\mathbb{R}^{2})$, $x^{u}(\cdot,\xi)=x(\cdot,\xi)$ solves
$\dot{x}(t,\xi)=\begin{bmatrix}0&-\xi_{1}\\ \xi_{1}&0\end{bmatrix}x(t,\xi)+\begin{bmatrix}u_{1}(t)\\ u_{2}(t)\end{bmatrix}+\begin{bmatrix}\xi_{2}\\ \xi_{3}\end{bmatrix},\quad t\in(0,1),\quad x(0,\xi)=\begin{bmatrix}1+\xi_{4}\\ \xi_{5}\end{bmatrix}$.
Moreover, the function $\psi(\cdot)$ is the sum of $(\alpha/2)\|\cdot\|_{L^{2}(0,1;\mathbb{R}^{2})}^{2}$ with $\alpha=10^{-3}$, and the indicator function of the set $\{u\in L^{2}(0,1;\mathbb{R}^{2})\colon\|u_{i}\|_{L^{\infty}(0,1;\mathbb{R})}\leq 3,\quad i=1,2\}$. The random variables $\xi_{1},\ldots,\xi_{5}$ are independent, and uniformly distributed on $0,2\pi]$, [5,5][-5,5], [5,5][-5,5], [1/2,1/2][-1/2,1/2], and [1/2,1/2][-1/2,1/2], respectively. The Cartesian product of these intervals defines the sample space $\Xi$.
Figure 1 depicts the nominal and reference solutions. Figure 2 depicts convergence rates for the SAA optimal values and critical points. These empirical rates are close to the theoretical rates established in sections 6 and 7. We have $\text{EPP}=4.14307$, $\text{RP}=3.85392$, yielding $\text{VSS}=0.28915$ and hence a relative VSS of $6.98\%$.
8.2 Vaccination scheduling for epidemic control under model parameter uncertainty
Figure 3: For the vaccination scheduling problem formulated in section 8.2, nominal solution (left) and reference solution (right).
We formulate an optimal control problem to obtain vaccination schedules for an epidemic disease under model parameter uncertainty. This problem formulation is inspired by the optimal control problem formulated in [15, sect. 4] (see also [12, Chap. 13]). The dynamics are based on an SEIR model described in [15, sect. 4.2], which tracks the number of individuals who are susceptible (S), exposed (E), infected (I), and recovered/vaccinated (R), as well as the total population. The model relies on six key parameters: natural death rate, disease death rate, birth rate, infection incidence, infection rate, and recovery rate. We account for uncertainty by treating these six parameters as random variables. Our control variable is the time-dependent vaccination rate for susceptible individuals. The goal is to minimize the expected number of infected individuals and the cost of vaccination.
We formulate the optimal control problem
$\min_{u\in L^{2}(0,t_{f};\mathbb{R})}\,(1/10)\mathbb{E}\Big[\int_{0}^{t_{f}}I^{u}(t,\xi)\mathrm{d}t\Big]+\psi(u),$ (19)
where for each control $u(\cdot)\in L^{2}(0,t_{f};\mathbb{R})$ with $0\leq u(t)\leq 0.9$ for a.e. $t\in(0,t_{f})$ and parameter $\xi\coloneqq(a,b,c,d,e,g)\in\Xi$, the states $S^{u}(\cdot,\xi)$, $E^{u}(\cdot,\xi)$, $I^{u}(\cdot,\xi)$, $R^{u}(\cdot,\xi)$, $M^{u}(\cdot,\xi)$ solve the SEIR model (see [15, eqns. (4.1)–(4.5)])
$\dot{S}(t)$ $=bM(t)-dS(t)-cS(t)I(t)-u(t)S(t),$ $S(0)$ $=S_{0},$
$\dot{E}(t)$ $=cS(t)I(t)-(e+d)E(t),$ $E(0)$ $=E_{0},$
$\dot{I}(t)$ $=eE(t)-(g+a+d)I(t),$ $I(0)$ $=I_{0},$
$\dot{R}(t)$ $=gI(t)-dR(t)+u(t)S(t),$ $R(0)$ $=R_{0},$
$\dot{M}(t)$ $=(b-d)M(t)-aI(t),$ $M(0)$ $=M_{0}.$
Following [15, p. 76], we choose $\psi(\cdot)$ as the sum of the indicator function of $\{u\in L^{2}(0,t_{f};\mathbb{R})\colon 0\leq u(t)\leq 0.9\text{ for a.e. }t\in(0,t_{f})\}$ and $(\alpha/2)\|\cdot\|_{L^{2}(0,t_{f};\mathbb{R})}^{2}$ with $\alpha=2$. For the numerical solution, we removed the fourth state equation, as none of the other states depend on $R(\cdot)$. The control problem (19) can be formulated as an instance of the risk-neutral problem (1) using a time transformation and incorporating an additional state variable.
We describe the parameter values and our choice of the random vector $\xi\in\mathbb{R}^{6}$ used for our numerical simulations. As in [15, Table 1], we use the initial states $S_{0}=1000$, $E_{0}=100$, $I_{0}=50$, and $R_{0}=15$. Let $M_{0}=S_{0}+E_{0}+I_{0}+R_{0}$, and let $t_{f}=20$. We choose the nominal parameter $\bar{\xi}=(0.2,0.525,0.001,0.5,0.5,0.1)$ for $\xi=(a,b,c,d,e,g)$ using the parameter values in [15, Table 1]. We construct the random variables $\xi_{j}$, $j=1,\ldots,6$, through random relative perturbations of the nominal parameter $\bar{\xi}$. Specifically, for $\sigma\coloneqq 0.15$ and independent random variables $\rho_{j}$, $j=1,\ldots,6$, each uniformly distributed on the interval [1,1][-1,1], we define
$\xi_{j}\coloneqq(1+\sigma\rho_{j})\bar{\xi}_{j},\quad j=1,\ldots,6.$
Therefore, $\Xi$ is chosen as the Cartesian product of the intervals $(1-\sigma)\bar{\xi}_{j},(1+\sigma)\bar{\xi}_{j}]$, $j=1,\ldots,6$.
Figure 3 depicts the nominal and reference solutions. Figure 4 depicts convergence rates for the SAA optimal values and critical points. These empirical rates closely match the theoretical rates established in sections 6 and 7. Given the potential nonconvexity of the vaccination scheduling problem, we cannot guarantee that the optimal values shown in Figure 4, as computed by the optimization solver, correspond to the true optimal values. We have $\text{EPP}=31.5168$, $\text{RP}=26.6276$, $\text{VSS}=4.88922$, and hence a relative VSS of $15.51\%$.
The optimal control problem (19) may not satisfy Assumption 2. Using the general invariance principle stated in [32, Thm. on p. 119] (see also the discussions in [8]), we can determine an invariant set for the SEIR dynamics that is independent of the feasible controls and parameters in $\Xi$. This result can be combined with the construction proposed in [28] to obtain modified dynamics and a corresponding optimal control problem that satisfy Assumption 2; see [28, eqns. (2.6) and (2.19), and Lemma 2.3] for further details.
Figure 4: For the vaccination scheduling problem formulated in section 8.2, convergence rate of the SAA optimal values as computed by the optimization solver (left) and of the reference criticality measure evaluated at SAA critical points (right). The empirical means $\widehat{\mathbb{E}}$ were computed using 5050 replications. The convergence rates were computed using least squares.
9 Discussion
We have considered risk-neutral optimal control of affine-control dynamics with random inputs. This manuscript has demonstrated nonasymptotic convergence rates for SAA optimal values and SAA critical points, which align with the typical Monte Carlo convergence rate. We have empirically validated these rates through numerical simulations. An open question remains whether the dependence on the strong convexity parameter in the convergence rates is optimal. This consideration may become particularly relevant for applications of our results to (statistical) inverse problems.
Acknowledgments
The second author thanks Ashwin Pananjady for insightful discussions on entropy integral bounds, optimality, and regularization. The authors thank the two anonymous referees for their comments and suggestions, which have improved the quality of this manuscript.
Reproducibility of computational results
Computer code that allows the reader to reproduce the computational results in this manuscript is available at https://doi.org/10.5281/zenodo.15529367.
Appendix A Optimality conditions
We state essentially known first-order necessary optimality conditions for composite optimization in Hilbert spaces. We use these optimality conditions to derive first-order optimality conditions for the control problem under uncertainty and its SAA problem, and to define criticality measures. These optimality conditions are used to justify the definition of certain criticality measures.
We consider the composite optimization problem
$\min_{u\in\mathrm{dom}(\psi)}\,G(u)+\psi(u).$ (20)
We formulate assumptions on the optimization problem (20).
Assumption 4.
  1. (a)
    The space $U$ is a real Hilbert space.
  2. (b)
    The function $\psi\colon U\to(-\infty,\infty]$ is proper, lower semicontinuous, and strongly convex with parameter $\alpha>0$.
  3. (c)
    The set $U_{1}\subset U$ is convex with $\mathrm{dom}(\psi)\subset U_{1}$. The function $G\colon U_{1}\to\mathbb{R}$ is Gateaux differentiable on $\mathrm{dom}(\psi)$ relative to $U_{1}$.
Let Assumption 4 hold true. We say that $\bar{u}\in\mathrm{dom}(\psi)$ is a critical point of (20) if $\nabla G(\bar{u})\in\partial\psi(\bar{u})$. Here, $\partial\phi(u)$ denotes the subdifferential of $\phi\colon U\to(-\infty,\infty]$ at $u\in U$. Let us define $\psi_{\alpha}(u)\coloneqq\psi(u)-(\alpha/2)\|u\|_{U}^{2}$.
The following essentially known fact characterizes the critical points of (20).
Lemma A.1.
Let Assumption 4 be satisfied. If $\bar{u}\in\mathrm{dom}(\psi)$ is a local solution to (20), then $\nabla G(\bar{u})\in\partial\psi(\bar{u})$. If $\bar{u}\in\mathrm{dom}(\psi)$, then $\bar{u}$ is a critical point of (20) if and only if
$\bar{u}=\mathrm{prox}_{\gamma\psi_{\alpha}}(\bar{u}-\gamma\nabla G(\bar{u})-\alpha\gamma\bar{u})\quad\text{for any and all}\quad\gamma>0.$
Proof.
The necessary first-order optimality condition is a direct consequence of well-known statements. Let us define $G_{\alpha}(u)\coloneqq G(u)+(\alpha/2)\|u\|_{U}^{2}$. We have $\partial\psi(u)=\partial\psi_{\alpha}(u)+\alpha u$. Therefore, $\bar{u}$ is a critical point for (20) if and only if $\nabla G_{\alpha}(\bar{u})\in\partial\psi_{\alpha}(\bar{u})$ if and only if $\gamma\nabla G_{\alpha}(\bar{u})\in\partial[\gamma\psi_{\alpha}(\bar{u})]$. Combined with [3, Prop. 12.26], we obtain the characterizations of critical points of (20). ∎
Appendix B Covering numbers of vector-valued functions
We provide upper bounds on the covering numbers of $\overline{\mathbb{B}}_{W_{1}^{2}(0,1;\mathbb{R}^{m})}(0;1)$ with respect to the $L^{2}(0,1;\mathbb{R}^{m})$-norm. The covering numbers may be nonoptimal with respect to the space dimension $m$, but are optimal with respect to the covering radii. In the main text, we use these covering numbers to establish those of a deterministic set containing all SAA critical points.
Proposition B.1.
For some $\varrho>0$ and each $m\in\mathbb{N}$, the binary logarithm of the $\nu$-covering number of $\overline{\mathbb{B}}_{W_{1}^{2}(0,1;\mathbb{R}^{m})}(0;1)$ with respect to the $L^{2}(0,1;\mathbb{R}^{m})$-norm does not exceed $\varrho\sqrt{m}(m/\nu)$ for all sufficiently small $\nu>0$.
Proof.
According to Theorem 1.7 in [5], the binary logarithm of the $\nu$-covering number of $\overline{\mathbb{B}}_{W_{1}^{2}(0,1)}(0;1)$ with respect to the $L^{2}(0,1)$-norm is proportional to $(1/\nu)$ for all sufficiently small $\nu>0$. In particular, there exists $\varrho>0$ such that for each $\nu>0$, this $\nu$-covering number does not exceed $2^{(\varrho/2)(1/\nu)}$. Fix $\nu>0$. Let $y_{1},\ldots,y_{K}$ be a $(\nu/\sqrt{m})$-cover of $\overline{\mathbb{B}}_{W_{1}^{2}(0,1)}(0;1)$ with respect to the $L^{2}(0,1)$-norm. We have $K\leq 2^{(\varrho/2)(\sqrt{m}/\nu)}$. We consider $\mathcal{W}_{K}\coloneqq\{w\in L^{2}(0,1;\mathbb{R}^{m})\ \colon w_{i}\in\{y_{1},\ldots,y_{K}\},\quad i=1,\ldots,m\}$. Fix $v\in\overline{\mathbb{B}}_{W_{1}^{2}(0,1;\mathbb{R}^{m})}(0;1)$. Hence $v_{i}\in\overline{\mathbb{B}}_{W_{1}^{2}(0,1)}(0;1)$, $i=1,\ldots,m$. We can choose $w\in\mathcal{W}_{K}$ such that for each $i\in\{1,\ldots,m\}$, $\|v_{i}-w_{i}\|_{L^{2}(0,1)}\leq\nu/\sqrt{m}$. Hence
$\|v-w\|_{L^{2}(0,1;\mathbb{R}^{m})}^{2}=\sum_{i=1}^{m}\|v_{i}-w_{i}\|_{L^{2}(0,1)}^{2}\leq\nu^{2}.$
The number of elements of $\mathcal{W}_{K}$ does not exceed $K^{m}\leq 2^{(\varrho/2)m(\sqrt{m}/\nu)}$.
Now, let $Y_{0}=\overline{\mathbb{B}}_{W_{1}^{2}(0,1;\mathbb{R}^{m})}(0;1)$ and $Y=L^{2}(0,1;\mathbb{R}^{m})$. From the proof of Lemma 8.2-2 (c) in [11], we deduce $\mathcal{N}(Y_{0};\nu)\leq\mathcal{N}(Y,Y_{0};\nu/2)$ for all $\nu>0$. Combining the pieces ensures the assertion. ∎
Appendix C Uniform expectation bounds in Hilbert spaces
We establish a uniform expectation bound for sample averages of independent Hilbert space-valued random variables using a chaining argument (see, for example, [7, 30]) and an exponential moment bound [22]. The uniform expectation bound allows us to establish convergence rates for SAA optimal values and SAA critical points, as demonstrated in the main text.
Throughout the section, let $(\Theta,\mathcal{A},\mathbb{P})$ be a complete probability space, and let $\xi$ be a $\Xi$-valued random element, where $\Xi$ is a complete separable metric space.
Theorem C.1.
Let $H$ be a real, separable Hilbert space, let $C\subset H$ be a nonempty, closed set with diameter $D>0$, and let $G\colon C\times\Xi\to H$ be a Carathéodory mapping with $\mathbb{E}[\|G(x,\xi)\|_{H}]<\infty$ and $\mathbb{E}[G(x,\xi)]=0$ for each $x\in C$. Suppose that there exists a constant $M>0$ such that
$\mathbb{E}[\cosh(\lambda\|G(x,\xi)-G(y,\xi)\|_{H})]\leq\mathrm{e}^{(1/2)M^{2}\lambda^{2}\|x-y\|_{H}^{2}}\;\text{for all}\;x,y\in C,\;\lambda\geq 0.$ (21)
Let $\xi^{1},\xi^{2},\ldots$ be independent $\Xi$-valued random elements defined on $(\Theta,\mathcal{A},\mathbb{P})$, each having the same distribution as $\xi$. We define $\widehat{G}_{N}(x)\coloneqq(1/N)\sum_{i=1}^{N}G(x,\xi^{i})$. Then for each $x_{0}\in C$ and $N\in\mathbb{N}$,
$\mathbb{E}[\sup_{x\in C}\,\|\widehat{G}_{N}(x)\|_{H}]\leq\mathbb{E}[\|\widehat{G}_{N}(x_{0})\|_{H}]+\frac{4\sqrt{3}M}{\sqrt{N}}\int_{0}^{D/2}\sqrt{\ln(2\mathcal{N}(C,\varepsilon))}\,\mathrm{d}\varepsilon.$
Under the hypotheses of Theorem C.1, we have (cf. [33, p. 79])
$\mathbb{E}[\|\widehat{G}_{N}(x_{0})\|_{H}]\leq\sqrt{\frac{\mathbb{E}[\|G(x_{0},\xi)\|_{H}^{2}]}{N}}.$
We leverage this bound in the main text.
Before we prepare our proof of Theorem C.1, we comment on the condition (21).
Remark C.2.
  1. (a)
    For a mean-zero random variable $Z\colon\Xi\to\mathbb{R}$, the condition, $\mathbb{E}[\cosh(\lambda|Z|)]\leq\exp(\lambda^{2}\tau^{2}/2)$ for all $\lambda\in\mathbb{R}$ and some $\tau>0$, is equivalent to $Z$ being sub-Gaussian (cf. [20, Lem. B.2]). This provides one motivation for using the condition (21) to model sub-Gaussian behavior.
  2. (b)
    If $G(\cdot,\xi)$ is Lipschitz continuous with Lipschitz constant $M$ for each $\xi\in\Xi$, then (21) holds true. This is a consequence of the inequality $\cosh(x)\leq\exp(x^{2}/2)$ valid for all $x\in\mathbb{R}$.
We prepare our proof of Theorem C.1. The following lemma provides a basic upper bound on the expected pointwise maximum of sub-Gaussian random variables.
Lemma C.3 (see [20, Lem. B.5]).
Let $\sigma>0$. If $Z_{k}:\Theta\to\mathbb{R}$ are random variables with $\mathbb{E}[\cosh(\lambda|Z_{k}|)]\leq\exp(\lambda^{2}\sigma^{2}/2)$ for all $\lambda\in\mathbb{R}$ and $k=1,2,\ldots,K$, then $\mathbb{E}[\max_{1\leq k\leq K}|Z_{k}|]\leq\sigma\sqrt{2\ln(2K)}$.
The following proposition is inspired by entropy integral bounds for sub-Gaussian processes (see, for example, [6, 7]).
Proposition C.4.
Let $(Y_{1},d_{Y_{1}})$ be a separable metric space, let $C\subset Y_{1}$ be a nonempty, closed set with diameter $D_{1}\in(0,\infty]$, let $Y_{2}$ be a separable Banach space, and let $G\colon C\times\Xi\to Y_{2}$ be a Carathéodory mapping. Suppose that
$\mathbb{E}[\cosh(\lambda\|G(x,\xi)-G(y,\xi)\|_{Y_{2}})]\leq\exp(\lambda^{2}d_{Y_{1}}(x,y)^{2}/2)\;\text{for all}\;x,y\in C,\;\lambda\geq 0.$ (22)
Then for all points $x_{0}\in C$, we have
$\mathbb{E}[\sup_{x\in C}\,\|G(x,\xi)\|_{Y_{2}}]\leq\mathbb{E}[\|G(x_{0},\xi)\|_{Y_{2}}]+4\sqrt{2}\int_{0}^{D_{1}/2}\sqrt{\ln(2\mathcal{N}(C,\varepsilon))}\,\mathrm{d}\varepsilon.$ (23)
Proof.
We adapt the proof of Theorem 3.1 in [6, p. 95]. If the right-hand side of (23) is infinite, then there is nothing to show. Next, let the right-hand side of (23) be finite. Hence $C$ is totally bounded and $D_{1}$ is finite. We define $\varepsilon_{0}\coloneqq D_{1}$ and $\varepsilon_{k}\coloneqq(1/2)^{k}\varepsilon_{0}$. For each $k\in\mathbb{N}$, let $C_{k}$ be a $\varepsilon_{k}$-net in the set $C$ with $\mathcal{N}(C,\varepsilon_{k})$ elements. We define $C_{\infty}\coloneqq\cup_{k=0}^{\infty}C_{k}$. The set $C_{0}$ contains only one point. We choose this point to be $x_{0}$.
Fix $k\in\mathbb{N}\cup\{0\}$. Following [6, p. 94], we define a mapping $\pi_{k}\colon C\to C_{k}$ such that $\pi_{k}(x)=x$ if $x\in C_{k}$ and $\pi_{k}(x)$ equals a point in $C_{k}$ closest to $x$ if $x\not\in C_{k}$.
Let $x\in C_{\infty}$. Then there exists $K(x)\in\mathbb{N}\cup\{0\}$ such that $x\in C_{K(x)}$. We define $x_{K(x)}\coloneqq x$, and $x_{k-1}\coloneqq\pi_{k-1}(x_{k})$, for $k=1,\ldots,K(x)$. We have
$G(x,\xi)=G(x_{0},\xi)+\sum_{k=1}^{K(x)}(G(x_{k},\xi)-G(x_{k-1},\xi)).$
Hence
$\|G(x,\xi)\|_{Y_{2}}\leq\|G(x_{0},\xi)\|_{Y_{2}}+\sum_{k=1}^{K(x)}\max_{x\in C_{k}}\|G(x,\xi)-G(\pi_{k-1}(x),\xi)\|_{Y_{2}}.$
Combined with (22), $d_{Y_{1}}(x,\pi_{k-1}(x))\leq\varepsilon_{k-1}$ for all $x\in C_{k}$, and Lemma C.3, we find that for $k=1,2,\ldots$,
$\mathbb{E}[\max_{x\in C_{k}}\,\|G(x,\xi)-G(\pi_{k-1}(x),\xi)\|_{Y_{2}}]\leq\varepsilon_{k-1}\sqrt{2\ln(2\mathcal{N}\big(C,\varepsilon_{k}\big))}.$
We obtain
$\mathbb{E}[\sup_{x\in C_{\infty}}\,\|G(x,\xi)\|_{Y_{2}}]$ $\leq\mathbb{E}[\|G(x_{0},\xi)\|_{Y_{2}}]+\sum_{k=1}^{\infty}\,\varepsilon_{k-1}\sqrt{2\ln(2\mathcal{N}\big(C,\varepsilon_{k}\big))}.$
Since the set $C_{\infty}$ is dense in $C$, $G$ is a Carathéodory mapping, and
$\varepsilon_{k-1}\sqrt{2\ln(2\mathcal{N}\big(C,\varepsilon_{k}\big))}\leq 4\int_{\varepsilon_{k+1}}^{\varepsilon_{k}}\sqrt{2\ln(2\mathcal{N}\big(C,\varepsilon\big))}\,\mathrm{d}\varepsilon,$
we obtain the assertion. ∎
The next lemma provides an exponential moment inequality. It is a direct consequence of Theorem 3 in [22] and Lemma 1 in [16], and has been established in [20].
Lemma C.5 (see [20, Lem. B.4]).
Let $\tau>0$, let $H$ be a real, separable Hilbert space, and let $Z_{i}:\Theta\to H$ be independent, mean-zero random vectors such that $\mathbb{E}[\cosh(\lambda\|Z_{i}\|_{H})]\leq\exp(\lambda^{2}\tau^{2}/2)$ for all $\lambda\geq 0$ and $i\in\{1,2,\ldots,N\}$. Then for each $\lambda\geq 0$,
$\mathbb{E}[\cosh(\lambda\|Z_{1}+\cdots+Z_{N}\|_{H})]\leq\exp(3\lambda^{2}\tau^{2}N/4).$
Now we are ready to establish Theorem C.1.
Proof of Theorem C.1.
Defining the metric $d_{Y_{1}}(x,y)\coloneqq\sqrt{3/2}(M/\sqrt{N})\|x-y\|_{H}$, Propositions C.4 and C.5 imply the assertion. ∎
The following example demonstrates an upper bound on the integral in (23) under a bound on the covering numbers. We use it in the main text to establish mean convergence rates for SAA optimal values and criticality measures.
Example C.6.
Let $c>0$ be a constant, and let $C\subset Y$ be a nonempty subset of a metric space $(Y,d_{Y})$ with diameter $D>0$. If $\mathcal{N}(C,\varepsilon)\leq 2^{c/\varepsilon}$ for all $\varepsilon\in(0,D]$, then
[…]